# Principled priors for Bayesian inference of circular models

### v2

@@ 1 Introduction
Advancements in computational power have provided researchers with access to massive and diverse datasets across various fields, enabling the exploration of complex scientific and practical challenges. However, tools for analyzing data with complex structures, such as directional or periodic characteristics, remain underdeveloped, and many of these datasets require specialized methods for analysis. For instance, wind directions and bird migration paths are naturally represented on a compass, while hospital patient arrival times and passenger density fluctuations follow a 24-hour clock (Mardia & Jupp, 2009). The orientation of earthquake epicenters, directions of cosmic rays and stellar objects in astronomy (Ley & Verdebout, 2017), joint angles prone to injury, typhoon trajectories, and the analysis of angular components in multivariate extreme value statistics are additional examples of data that align with circular scales. These examples underscore the growing need for statistical frameworks capable of handling circular data.
@@
Directional statistics offers appropriate tools for analyzing such data, with circular distributions — probability distributions defined on the circumference of a circle (Jammalamadaka, 2001) and characterized by angular measures or radians — playing a central role. These methods are crucial for modeling data in fields such as meteorology, earth sciences, bioinformatics, ecology, medicine, genetics, neurology, astronomy, and beyond (Ley & Verdebout, 2017; Pewsey & García-Portugués, 2021).
@@
Applying Bayesian methods to circular data, however, poses significant challenges, particularly in the selection of priors — a pivotal step in Bayesian analysis. Unlike Euclidean distributions, circular distributions have unique parameterizations and behaviors that are less well-understood, often requiring specialized approaches. For example, in Wallace & Dowe (1993), priors for the concentration parameter $\kappa$ of von Mises distribution (Section 2.1) are constructed through the techniques of Minimum Message Length. These priors are also used in Dowe et al. (1996) and Marrelec & Giron (2024). Another popular prior for von Mises distribution is the joint conjugate prior proposed in Damien & Walker (1999). In addition, in the general procedure for Bayesian analysis with wrapped distributions proposed by Ravindran & Ghosh (2011), the $\operatorname{Beta}\left(a,a\right)$ prior is employed. Moreover, many studies try to avoid the issues in constructing Bayesian circular models. For example, Nuñez-Antonio et al. (2011) proposed to fit the Bayesian circular model through the projected normal distribution with normal priors.
Limited theoretical development in this area further complicates prior selection, increasing the risk of poorly informed choices that can either dominate the posterior or lead to poorly performing models. It is essential therefore to examine the necessary conditions which can cultivate proper priors so as to prevent overfitting, especially when little prior knowledge about the parameters is available.
@@
To address these challenges, two goals are set in this research. First, efforts are made to establish a principled, practical and systematic framework for circular prior selection to account for the unique complexities of circular distributions including base model selection and distance computation. This framework is inspired by the Penalized Complexity (PC) prior framework proposed by Simpson et al. (2017). The PC prior framework is a reliable choice for constructing default priors, which can balance the model complexity and prior informativeness, particularly in scenarios with limited prior knowledge, when objective or uninformative priors are usually considered. To develop proper priors in circular distribution scenarios, we need to adapt the framework components to the new circumstances to help derive proper priors for this new situation so as to check whether the PC priors are optimal ones. Second, with the guidance of the framework, explicit expressions for Penalized Complexity (PC) priors for the most commonly used circular distributions are to be developed so as to tackle the methodological challenges inherent in these distributions. Moreover, in-depth comparisons are conducted between PC priors and common priors found in literature, together with extensive simulation studies to evaluate their performance and demonstrate their practical relevance through real-world data applications.
Our efforts of adapting the framework for circular prior selection together with the developed PC priors will provide a proper foundation for reliable Bayesian analysis of circular data and future applications.
@@
The structure of this paper is as follows: Section 2 reviews the most commonly used circular distributions, their properties, and existing priors in the literature. Section 3 introduces the proposed framework for prior selection and the procedure for deriving PC prior for circular distributions. Section 4 derives explicit formulations for widely used circular distributions. Section Section 5 evaluates the proposed priors through comparison studies and simulation studies, while Section Section 6 demonstrates their application to real-world datasets. Finally, Section Section 7 discusses broader implications and potential directions for future research.
@@ 2 Preliminaries
One of the most important characteristics of circular data is that the data are defined on $\left[0,2\pi\right)$ instead of on $\mathbb{R}$ and a proper circular distribution should satisfy $\lim_{x\rightarrow 0_{+}}\pi\left(x\right)=\lim_{x\rightarrow 2\pi_{-}}\pi\left(x\right)$. A variety of circular distributions have been proposed to model circular data with this feature. Apart from the simplest, the circular uniform distribution (Equation 1), most circular distributions are constructed by using one of the following four general approaches: wrapping, conditioning, projection, and perturbation (Ley & Verdebout, 2017). Notably, many of these distributions include the circular uniform distribution as a special case (Figure 3), defined by a probability density function as follows:
$\pi_{\mathcal{U}}\left(x\right)$ $=\frac{1}{2\pi},\quad x\in\left[0,2\pi\right).$ (1)
Three of the most widely used circular distributions - the von Mises (vM) distribution, the cardioid distribution, and the wrapped Cauchy (WC) distribution — as highlighted by Ley & Verdebout (2017), serve as the foundation for many generalizations and extensions in circular statistics, such as the Jones-Pewsey distribution (Jones & Pewsey, 2005) and the Kato-Jones distribution (Kato & Jones, 2010), underscoring their importance in the field.
To gain a better understanding of the properties of these circular distributions and what priors are set in them, a brief discussion is given below together with a figure (Figure 3) to illustrate the relationship among these distributions.
@@ 2.1 Von Mises Distribution
The von Mises distribution (von Mises, 1918), often referred to as the circular normal distribution (Gumbel et al., 1953), is the circular analogue of the normal distribution. It is one of the most widely used and versatile circular distributions (Mardia & Jupp, 2009). Its probability density function is given by:
$\pi_{\mathcal{VM}}\left(x;\mu,\kappa\right)$ $=\frac{1}{2\pi\mathcal{I}_{0}\left(\kappa\right)}\exp\left\{\kappa\cos\left(x-\mu\right)\right\},\quad\mu\in\left[0,2\pi\right),\quad\kappa\in\left[0,\infty\right),$ (2)
where $\mu$ is the location parameter, $\kappa$ is the concentration parameter, and $\mathcal{I}_{a}\left(\cdot\right)$ is the modified Bessel function of the first kind of order $a\in\mathbb{N}$. Figure 1 illustrates the behavior of the density with different $\kappa$ values when $\mu=\pi$.
From the density function and Figure 1, it is clear that the von Mises distribution includes the circular uniform distribution as a special case when $\kappa=0$. In this scenario, there is no concentration in the data, while $\kappa\to\infty$ indicates that the data are highly concentrated around the mean direction $\mu$. Thus, $\kappa$ can be interpreted as analogous to the precision of the normal distribution ($1/\sigma^{2}$). Since the Gamma distribution is a common prior choice for the precision in normal distributions, the $\operatorname{Gamma}(a,b)$ distribution is similarly well-suited as a prior for $\kappa$.
For the location parameter $\mu$, the circular uniform prior is often employed in the absence of prior information. Unlike location parameters in distributions on $\mathbb{R}$, this choice is less problematic for circular distributions since their variables are bounded within $\left[0,2\pi\right)$. As a result, the circular uniform prior is less prone to lead to a posterior with heavier tails or substantial bias, which could occur for location parameters in unbounded domains.
Dowe et al. (1996) and Marrelec & Giron (2024) both adopted a circular uniform prior for $\mu$ and considered two different priors for $\kappa$ originally proposed by Wallace & Dowe (1993). These priors are defined as:
@@
Another family of priors used is the conjugate priors. Guttorp & Lockhart (1988) proposed a joint conjugate prior for $\mu$ and $\kappa$, expressed as:
$\pi\left(\mu,\kappa\right)$ $=\frac{1}{2\pi\mathcal{I}_{0}\left(\kappa R_{0}\right)}\exp\left\{\kappa R_{0}\cos\left(x-\mu_{0}\right)\right\}=\propto\left[\mathcal{I}_{0}\left(\kappa\right)\right]^{-c}\exp\left\{\kappa R_{0}\cos\left(x-\mu_{0}\right)\right\}$ (4)
where $c$, $R_{0}$ and $\mu_{0}$ are prior hyperparameters. When $c\in\mathbb{N}$, $c$ can be interpreted as representing $c$ prior observations concentrated around the direction $\mu_{0}$, with $R_{0}$ corresponding to the component of the resultant vector in the known direction (Damien & Walker, 1999).
Despite its intuitive interpretation, deriving the posterior distribution using this prior introduces three additional hyperparameters and several latent variables during the sampling process (see Damien & Walker 1999 for details). This process, however, increases computational complexity and leads to greater variation in the posterior estimates.
Figure 1: Density of von Mises distributions with low (left) and high (right) concentration.
@@ 2.2 Cardioid Distribution
Another classic circular distribution is the cardioid distribution (Figure 2, left) introduced by Jeffreys (1948, P.302). It is a cardioid perturbation of the circular uniform distribution (Ley & Verdebout, 2017) with probability density function given by
$\pi_{\mathcal{C}}\left(x;\mu,\ell\right)$ $=\frac{1}{2\pi}\left(1+2\ell\cos\left(x-\mu\right)\right),\quad\mu\in\left[0,2\pi\right),\quad\ell\in\left[0,1/2\right).$ (5)
Here, $\ell$ is the concentration parameter which controls the deviation from the circular uniform distribution. Larger values of $\ell$ result in stronger departures from uniformity. This distribution is primarily used as a small-concentration approximation to the von Mises distribution, as noted by Mardia & Jupp (2009).
Figure 2: Density of cardioid (left) and wrapped Cauchy (right) distributions.
This distribution is rarely used today and there is insufficient literature for setting priors for it. However, two reasonable choices arise for $\ell$ in this distribution which are the $\operatorname{Uniform}\left(0,0.5\right)$ prior and the $2\times\operatorname{Beta}\left(a,b\right)$ prior, since both priors fit the requirement of support being $\left[0,0.5\right)$.
@@ 2.3 Wrapped Cauchy Distribution
The wrapped Cauchy distribution (WC) (Figure 2, right), defined in Equation 6, is constructed by wrapping the Cauchy probability density around a unit circle (modulus $2\pi$). An advantage of the WC distribution is that it is a wrapped distribution with probability density function having a closed form (Ley & Verdebout, 2017). Further details of the WC distribution are discussed in Chapter 3.5.7 of Mardia & Jupp (2009). The probability density function is as follows:
$\pi_{\mathcal{WC}}\left(x;\mu,\rho\right)=\frac{1}{2\pi}\frac{1-\rho^{2}}{1+\rho^{2}-2\rho\cos\left(x-\mu\right)},\quad\text{where}\quad\mu\in[0,2\pi),\quad\rho\in\left[0,1\right).$ (6)
The parameter $\rho$ serves as a measure of concentration, with $\rho=0$ corresponding to the circular uniform distribution, indicating no concentration. As $\rho\to 1$, the data become increasingly concentrated, converging to a point mass distribution.
In practice, the WC distribution is primarily used to model data with high dispersion, and it is generally not the preferred choice for strongly concentrated data. Compared to the von Mises (vM) distribution (Figure 1), the WC distribution typically exhibits greater dispersion, making it better suited for applications with widely-spread data. For strongly concentrated data, the von Mises distribution offers a more accurate representation due to its robust theoretical properties and consistent behavior.
For Bayesian analysis with a WC distribution, one reasonable choice for the prior for $\rho\in\left[0,1\right]$ is a $\operatorname{Beta}\left(a,b\right)$ prior, which has support in $\left[0,1\right]$. Ravindran & Ghosh (2011) mentioned that $\operatorname{Beta}\left(a,a\right)$ is a class of non-informative prior for $\rho$. However, the beta prior behaves markedly differently with different choices for $a$ and $b$, and therefore they need to be chosen carefully.
Figure 3: Relationship between popular circular distributions.
The circular distributions and priors discussed above highlight the diversity and utility of existing methods for modeling circular data. However, many of the priors currently in use are either heuristic or tailored to specific cases, lacking a unified framework for selection and hyperparameter determination. This motivates the need for a principled and systematic approach for prior selection which can account for the unique properties of circular distributions. In the next section, we will introduce a principled framework for prior selection for circular distributions, and provide the procedure for constructing the PC prior in circular scenarios, offering a consistent methodology for addressing the challenges discussed here.

### v3

@@ 1 Introduction
Advancements in computational power have provided researchers with access to massive and diverse datasets across various fields, enabling the exploration of complex scientific and practical challenges. However, tools for analyzing data with complex structures, such as directional or periodic characteristics, remain underdeveloped, and many of these datasets require specialized methods for analysis. For instance, wind directions and bird migration paths are naturally represented on a compass, while hospital patient arrival times and passenger density fluctuations follow a 24-hour clock (Mardia and Jupp, 2009). The orientation of earthquake epicenters, directions of cosmic rays and stellar objects in astronomy (Ley and Verdebout, 2017; Pewsey and García-Portugués, 2021; Cabella and Marinucci, 2009), joint angles prone to injury, protein structure with angular measures in bioinformatics (Mardia et al., 2018; Boomsma et al., 2008), typhoon trajectories, and the analysis of angular components in multivariate extreme value statistics are additional examples of data that align with circular scales. These examples underscore the growing need for statistical frameworks capable of handling circular data.
@@
Directional statistics offers appropriate tools for analyzing such data, with circular distributions — probability distributions defined on the circumference of a circle (Jammalamadaka, 2001) and characterized by angular measures or radians — playing a central role. These methods are crucial for modeling data in fields such as meteorology, earth sciences, bioinformatics, ecology, medicine (Vuollo et al., 2016; Pardo et al., 2016), genetics, neurology, astronomy (Cabella and Marinucci, 2009; Marinucci and Peccati, 2011), image analysis (Jung et al., 2011; Esteves et al., 2018), text mining (Dhillon and Modha, 2001; Banerjee et al., 2005), machine learning (Sra, 2016), and beyond (Ley and Verdebout, 2017; Pewsey and García-Portugués, 2021).
@@
Applying Bayesian methods to circular data, however, poses significant challenges, particularly in the selection of priors — a pivotal step in Bayesian analysis. Unlike Euclidean distributions, circular distributions have unique parameterizations and behaviors, often requiring specialized approaches. For example, in Wallace and Dowe (1993), priors for the concentration parameter $\kappa$ of von Mises distribution are constructed through the techniques of Minimum Message Length (Section 2.1). These priors are also used in Dowe et al. (1996) and Marrelec and Giron (2024). Another popular prior for von Mises distribution is the joint conjugate prior proposed in Damien and Walker (1999). In addition, in the general procedure for Bayesian analysis with wrapped distributions proposed by Ravindran and Ghosh (2011), the $\operatorname{Beta}\left(a,a\right)$ prior is employed. Nuñez-Antonio et al. (2011) proposed to fit the Bayesian circular model through the projected normal distribution with normal priors. A prior for the location parameter is often more intuitive to formulate, while priors for the hyperparameters like dispersion parameters are notoriously hard to conceptualize. Moreover, when specific values of the hyperparamaters result in model complexity reduction, care should be taken in the prior construction.
Since circular variables are defined on a compact and curved space rather than in linear Euclidean space, their geometric properties make it difficult to develop strong intuition about their behavior. This further complicates prior selection, increasing the risk of using priors that either dominate the posterior, or possible lead to poorly performing models. Poorly chosen priors can cause overfitting in the Bayesian model. It is worth noting that overfitting in this article refers to the Informal Definition 1 of Simpson et al. (2017), which defines a prior $p\left(\xi\right)$ as forcing overfitting if it allocates insufficient mass to the simpler model, making it difficult for the posterior to recover a simpler model when supported by the data. A detailed discussion of the simpler model is presented in Section 3.1.
@@
To address these challenges, we set two goals for this research. Firstly, we establish a principled framework for the hyperparameters of a general circular model, based on a simpler circular model. This framework is inspired by the Penalized Complexity (PC) prior framework proposed by Simpson et al. (2017). The PC prior framework is a reliable choice for constructing default priors, which can balance model complexity and prior informativeness, particularly in scenarios with limited prior knowledge, when objective or uninformative priors are usually considered. Second, we derive the explicit expressions for the Penalized Complexity (PC) priors for the most commonly used circular distributions’ hyperparameters and provide a way to quantify prior information through a user-defined parameter.
@@
The structure of this paper is as follows: Section 2 reviews the most commonly used circular distributions, their properties, and existing priors in the literature. Section 3 introduces the proposed framework for prior selection and the procedure for deriving PC prior for circular distributions. Section 3.2 derives explicit formulations for widely used circular distributions. Section 4 evaluates the proposed priors through comparison studies and simulation studies, while Section 5 demonstrates their application to real-world datasets. Finally, Section 6 discusses broader implications and potential directions for future research.
@@ 2 Preliminaries
A proper circular distribution should have a probability density function $p\left(x\mid\xi\right)$ that satisfies $p\left(x\mid\boldsymbol{\xi}\right)=p\left(x+2\pi k\mid\boldsymbol{\xi}\right)$ for any integer $k$ and $x\in[0,2\pi)$, where $\boldsymbol{\xi}$ is the parameters for the probability density function. Most circular distributions are constructed by using one of the following four general approaches: wrapping, conditioning, projection, and perturbation (Ley and Verdebout, 2017).
Circular distributions in the wrapped family are constructed by taking a distribution defined on $\mathbb{R}$ and wrapping it around a circle (modulus by $2\pi$, e.g. wrapped Normal distribution, wrapped Cauchy distribution, wrapped double exponential distribution); the conditioning approach obtains circular distributions through constructing joint distribution of polar coordinates (radius $r$ and angle $\theta$) and finding the conditional distribution $p\left(\theta\mid r\right)$ through restricting the radius $r=1$ (e.g. von Mises distribution); the projection approach projects an distribution on $\mathbb{R}^{2}$ onto the unit circle (e.g. rojected Normal distribution); The perturbation approach provides flexible choices to extend a circular density to a more general form through multiplying it with a proper function (e.g. cardioid distribution). Details can be found in Chapter 2.2 of Ley and Verdebout (2017).
Notably, many of these distributions include the circular uniform distribution as a special case (Figure 1), defined by a probability density function given by $p_{\mathcal{U}}\left(x\right)=\frac{1}{2\pi},\quad x\in\left[0,2\pi\right)$.
Figure 1: Relationship between popular circular distributions. $\kappa$, $\ell$ and $\rho$ are the concentration parameters for von Mises, cardioid and wrapped Cauchy distributions.
Three of the most widely used circular distributions — the von Mises (vM) distribution, the cardioid distribution, and the wrapped Cauchy (WC) distribution — serve as the foundation for many generalizations and extensions in circular statistics (Ley and Verdebout, 2017), such as the Jones-Pewsey distribution (Jones and Pewsey, 2005) and the Kato-Jones distribution (Kato and Jones, 2010), underscoring their importance in the field. One commonality between these three distributions is that they all include two parameters: one location parameter and one concentration parameter. The concentration parameter is essentially a scaling parameter for circular distributions. When data is distributed on a unit circle (or sphere/hypersphere), the concept of the standard deviation, as defined for linear data, becomes meaningless, since its interpretation is no longer clear, particularly when the standard deviation exceeds half a circle ($\pi$). To describe the spread or sparsity of circular data, most circular distributions introduce a “concentration” parameter, which quantifies the extent of concentration (or dispersion) within the data, indicating how densely the data are clustered around a mean direction (or how widely they are spread). In this work, we focus on the von Mises (vM) distribution, the cardioid distribution, and the wrapped Cauchy (WC) distribution, and on their Bayesian inference.
@@ 2.1 Von Mises Distribution
The von Mises distribution (Mardia and Jupp, 2009), often referred to as the circular normal distribution (Gumbel et al., 1953), is the circular analogue of the normal distribution. It is one of the most widely used and versatile circular distributions (Mardia and Jupp, 2009). Its probability density function is given by:
$p_{\mathcal{VM}}\left(x\mid\mu,\kappa\right)$ $=\frac{1}{2\pi\mathcal{I}_{0}\left(\kappa\right)}\exp\left\{\kappa\cos\left(x-\mu\right)\right\},\quad\mu\in\left[0,2\pi\right),\quad\kappa\in\left[0,\infty\right),$ (1)
where $\mu$ is the location parameter, $\mathcal{I}_{a}\left(\cdot\right)$ is the modified Bessel function of the first kind of order $a\in\mathbb{N}$, and $\kappa$ is the concentration parameter. The plot for this density is given in Figure 2.
Figure 2: vM density for small (left) and large (right) $\kappa$ values with $\mu=\pi$ .
The von Mises distribution includes the circular uniform distribution as a special case when $\kappa=0$, while $\kappa\to\infty$ indicates that the data are highly concentrated around the mean direction $\mu$. Thus, $\kappa$ can be interpreted as analogous to the precision of the normal distribution ($1/\sigma^{2}$). Since the Gamma distribution is a common prior choice for the precision in normal distributions, the $\operatorname{Gamma}\left(a,b\right)$ distribution with density given in Equation 2 is naively considered as a prior for $\kappa$.
$p\left(x\mid a,b\right)$ $=\frac{b^{a}}{\Gamma\left(a\right)}x^{a-1}\exp\left\{-bx\right\}$ (2)
Guttorp and Lockhart (1988) proposed a joint conjugate prior for $\mu$ and $\kappa$, expressed as:
$p\left(\mu,\kappa\right)$ $\propto\left[\mathcal{I}_{0}\left(\kappa\right)\right]^{-c}\exp\left\{\kappa R_{0}\cos\left(x-\mu_{0}\right)\right\}$ (3)
where $c$, $R_{0}$ and $\mu_{0}$ are prior hyperparameters. When $\in\mathbb{N}, cc$ can be interpreted as representing $c$ prior observations concentrated around the direction $\mu_{0}$, with $R_{0}$ corresponding to the component of the resultant vector in the known direction (Damien and Walker, 1999).
Despite its intuitive interpretation, deriving the posterior distribution under this prior requires introducing additional hyperparameters and several latent variables during the sampling process (see Damien and Walker 1999 for details). This construction, however, increases computational complexity and often results in greater Monte Carlo variability of the posterior estimates, owing to the mixing and dependence among latent components. In some cases, particularly when $c$ and $R_{0}$ are not well chosen, or when their combinations convey weak prior information, it may also lead to wider posterior credible intervals.
Dowe et al. (1996) and Marrelec and Giron (2024) both considered two different priors for $\kappa$ originally proposed by Wallace and Dowe (1993) derived from minimum message length (MML). These priors are defined as:
@@
Minimum Message Length (MML) is a Bayesian information-theoretic restatement of Occam’s razor, favoring models that describe the data most efficiently. It encodes both the model and the data as one message, where a more complex model requires a longer description and is only preferred if it leads to a shorter overall message. In this sense, priors derived from MML, such as $h_{2}\left(\kappa\right)$ and $h_{3}\left(\kappa\right)$, naturally penalize overly complex or highly concentrated models unless the data provide strong evidence to support them.
@@ 2.2 Cardioid Distribution
The cardioid distribution a cardioid perturbation of the circular uniform distribution (Ley and Verdebout, 2017) with probability density function given by
$p_{\mathcal{C}}\left(x\mid\mu,\ell\right)$ $=\frac{1}{2\pi}\left(1+2\ell\cos\left(x-\mu\right)\right),\quad\mu\in\left[0,2\pi\right),\quad\ell\in\left[0,1/2\right).$ (5)
Here, $\ell$ is the concentration parameter which controls the deviation from the circular uniform distribution. Larger values of $\ell$ result in stronger departures from uniformity (see the left-hand-side plot in Figure 3). It is obvious from the plot that the cardioid density is much less concentrated compared with the von Mises density, even when the concentration parameter $\ell$ for the former density approaches its supremum. This property enable the cardioid distribution to be more appropriate for dispersed data (heavy-tailed data) compared with von Mises distribution. Thus, the cardioid distribution is primarily used as a small concentration approximation to the von Mises distribution, as noted by Mardia and Jupp (2009).
Figure 3: Cardioid density (left) and wrapped Cauchy density (right) with $\mu=\pi$ for different concentration parameter values.
Two reasonable prior choices for $\ell$ could be the $\operatorname{Uniform}\left(0,0.5\right)$ prior and the $0.5\times\operatorname{Beta}\left(a,b\right)$ prior, since both priors fit the requirement of support being $\left[0,0.5\right)$.
@@ 2.3 Wrapped Cauchy Distribution
The wrapped Cauchy distribution is one of the rare wrapped distribution that has an analytic density function, since its probability density can be expressed in closed form (Ley and Verdebout, 2017). Further details of the WC distribution are discussed in Chapter 3.5.7 of Mardia and Jupp (2009). The probability density function is given by:
$p_{\mathcal{WC}}\left(x\mid\mu,\rho\right)=\frac{1}{2\pi}\frac{1-\rho^{2}}{1+\rho^{2}-2\rho\cos\left(x-\mu\right)},\quad\text{where}\quad\mu\in[0,2\pi),\quad\rho\in\left[0,1\right).$ (6)
The parameter $\rho$ serves as a measure of concentration, with $\rho=0$ corresponding to the circular uniform distribution, indicating no concentration and as $\rho\to 1$, the data become increasingly concentrated, converging to a point mass distribution. The density plot for different $\rho$ values is given in the right-hand side of Figure 3.
For Bayesian analysis with a WC distribution, one reasonable choice for the prior for $\rho\in\left[0,1\right]$ is a $\operatorname{Beta}\left(a,b\right)$ prior, which has support in $\left[0,1\right]$. Ravindran and Ghosh (2011) mentioned that $\operatorname{Beta}\left(a,a\right)$ is a class of non-informative priors for $\rho$. However, the beta prior behaves markedly differently with different choices for $a$ and $b$, and therefore they need to be chosen carefully.
The circular distributions and priors discussed above highlight the diversity and utility of existing methods for modeling circular data. However, many of the priors currently in use are either heuristic or tailored to specific cases, lacking a unified framework. This motivates the need for a principled approach for prior selection which can account for the unique properties of circular distributions, which we present in the next section.